# Shen2018S18

### v1

@@ 4.1 Wide BH companions are overrepresented at low metallicity
Figure 4: Metallicity distribution of red giants within 1 kpc of the Sun, calculated from the Gaia XP metallicity catalog of Andrae et al. (2023, black) and APOGEE DR17 (Abdurro’uf et al., 2022, red). The metallicity of Gaia BH3 is marked with a shaded region: $\rm[M/H]=-1.82$ according to Andrae et al. (2023), and $\rm[M/H]=-2.2$ according to Gaia Collaboration et al. (2024). The fraction of giants with metallicities as low as Gaia BH1 is only $\sim 0.0001$.
@@
Both the XP and APOGEE samples suggest that only 1 in $\sim 10^{4}$ giants within 1 kpc have $\rm[M/H]<-2.0$. This very small fraction strongly implies that wide BH companions are overrepresented at low metallicity. If they were not, we could expect to find $\sim 10^{4}$ BH companions – just to red giants – within 1 kpc. The true number of BH + giant binaries within 1 kpc is unknown, but 0 have been discovered so far: the nearest such system is Gaia BH2 at $d=1.16$ kpc (El-Badry et al., 2023b). It seems quite unlikely that $10^{4}$ remain to be discovered (although published Gaia data are not yet sensitive to orbital periods as long as that of Gaia BH3), and so a more plausible explanation is that wide BH companions are more common at low metallicity.
@@ 4.2 More low-metallicity BHs are expected in DR4
There is also reason to believe that more BHs orbited by metal-poor stars exist in the solar neighborhood. Such stars are uniformly $\gtrsim 12$ Gyr old. This means that giants sample a narrow mass range of roughly $(0.78-0.80)\,M_{\odot}$. For a Kroupa IMF, there are $\sim$165 stars with $M=(0.1-0.78)\,M_{\odot}$ for every star with $M=(0.78-0.80)\,M_{\odot}$, suggesting that even within 1 kpc, other massive BHs are likely orbited by lower-mass metal-poor stars that are still on the main sequence. The mass distribution of companions to BHs of course may be different from the IMF, but it seems unlikely that $0.8\,M_{\odot}$ companions would be significantly more common than companions with mass of $0.7\,M_{\odot}$ or $0.6\,M_{\odot}$. At the distance and extinction of Gaia BH3, such companions would have $G<14$ for $M>0.75\,M_{\odot}$, or $G<16$ for $M>0.61\,M_{\odot}$. These magnitudes are too faint for Gaia to have measured multi-epoch RVs from RVS spectra, but bright enough that the astrometric orbit would be resolved with $\rm SNR\gtrsim 100$. With careful processing of the astrometric data, a source like Gaia BH3 could likely be detected even at $G=18-19$ – where individual astrometric measurements have uncertainties of 1-3 mas in the along-scan direction, a factor of 10-30 smaller than Gaia BH3’s photocenter ellipse (Lindegren et al., 2018) – although spurious solutions will be more abundant at low SNR. Gaia BH3-like systems can thus quite plausibly be detected with significantly lower-mass secondaries, and to larger distances, than Gaia BH3 itself.
@@ 5.1 Formation from an isolated binary
There are at least three straightforward channels to form binaries lie Gaia BH3 through isolated binary evolution:
  1. 1.
    The progenitor of the BH could have become a red supergiant but never overflowed its Roche lobe. The MESA models described in Section 3 have maximum radii of $1000-1800\,R_{\odot}$. Meanwhile, Monte Carlo kick simulations (Figure 1) suggest a median pre-SN separation of $\sim 4500\,R_{\odot}$, such that the BH progenitor would only have overflowed its Roche lobe if its radius exceed $\sim 3000\,R_{\odot}$. It thus seems quite possible that the BH progenitor and companion star never interacted.
  2. 2.
    The BH progenitor could have overflowed its Roche lobe shortly before core collapse, such that there was no time for the companion star to spiral very far inward. This does not necessarily require fine-tuning: the models predict that at the very low metallicities relevant to Gaia BH3, massive stars spend most of their post-main sequence evolution as blue/yellow supergiants that are too small ($R\lesssim 100\,R_{\odot}$) to overflow their Roche lobes for any plausible pre-SN orbit in Gaia BH3. They do eventually expand to become red supergiants, but only in the last $\lesssim 1000$ years before core collapse (Figure 3). Roche lobe overflow during this phase may not have left enough time for the companion star to spiral inward significantly.
  3. 3.
    The BH could have formed via chemically homogeneous evolution of a $\sim 35\,M_{\odot}$ star with initial rotation rate $\Omega/\Omega_{\rm crit}\gtrsim 0.45$. Such a star would never have become a red supergiant and so would have comfortably fit within its Roche lobe for any possible initial orbit.
Some caution is advisable in interpreting the maximum radii predicted by the MESA models. The radii of red supergiants are sensitive to a variety of uncertain inputs, such as the convective mixing length and the treatment of overshooting and semiconvection (Chun et al., 2018; Schootemeijer et al., 2019; Goldberg & Bildsten, 2020; Klencki et al., 2021). And no one has ever observed a massive star with a metallicity as low as Gaia BH3, so the evolutionary models are untested. However, I emphasize that (a) the modeling choices produce red supergiant temperatures and luminosities in reasonably good agreement with observations of massive stars in the LMC and SMC (e.g. Schootemeijer et al., 2019), which are the lowest-metallicity galaxies current observable with a large population of evolved massive stars, and (b) the radii predicted by these models are not unusually small compared to the predictions of other work (e.g., the references above). The orbit of Gaia BH3 is simply wide, and can accommodate even very large BH progenitors.
@@ 5.2 Dynamical formation
The fact that Gaia BH3’s orbit is consistent with having formed through isolated binary evolution does not, of course, rule out the possibility that it formed through dynamic interactions in a dense cluster (e.g. Rastello et al., 2023; Di Carlo et al., 2023; Tanikawa et al., 2024b). Indeed, Balbinot et al. (2024) report a likely association between Gaia BH3 and the ED-2 halo stellar stream, which was most likely formed from a low-mass metal-poor star cluster. This makes it quite natural to consider dynamical formation channels, and cluster models tuned to the progenitor of ED-2 will be able predict the population of BH + stellar binaries they should form (e.g. Kremer et al., 2018; Gieles et al., 2021). Association with a cluster does not, however, guarantee that Gaia BH3 formed dynamically: most massive stars form in clusters, and clusters contain both primordial and dynamically-formed binaries. The key uncertainties are the initial density and mass of the cluster and how rapidly it dissolved. If Gaia BH3 formed in a cluster, this would require even weaker kicks than implied by Figure 1, because the escape velocity of plausible low-mass clusters is only a few $\rm km\,s^{-1}$.
@@
Gaia BH3 is just at the separation beyond which BH+star binaries can avoid a common envelope. Irrespective of whether Gaia BH3 in particular formed from an isolated binary, the models presented here imply that isolated binary evolution produces binaries with orbits similar to Gaia BH3 – which is not the case for closer binaries like Gaia BH1 and BH2 – and future Gaia data releases are likely to find these binaries.
@@
Constraints on the separation distribution of BH binaries from DR4 will provide some insight into their dominant formation channel. If these systems form mainly by isolated binary evolution, models predict the BH + low mass star period distribution to rise steeply at $P_{\rm orb}\gtrsim 10$ yr (e.g. Breivik et al., 2017; Chawla et al., 2022), the period beyond which a significant fraction of binaries will have avoided a common envelope inspiral. For dynamical formation channels, the period distribution will depend on the mass of the clusters in which most systems are formed, but there is little reason to expect a steep increase at $P_{\rm orb}\gtrsim 10$ yr. Fortunately, Gaia will obtain an 11-year observational baseline, enabling robust orbital constraints at periods up to $\sim 20$ years, and perhaps even significantly longer (e.g. Andrews et al., 2023).
@@
The metallicity distribution of BH binaries will also help constrain how they formed. The models presented here suggest that isolated binary evolution is more promising at low metallicity, because the compact nature of massive low-metallicity stars means that they can avoid overflowing their Roche lobes and still form BH + low-mass star binaries in the separation range probed by Gaia DR4. On the other hand, the vast majority of dense clusters that have contributed stars to the solar neighborhood had metallicities near solar. If dynamical interactions form most wide BH binaries, most systems should likely have metallicities near solar, similar to Gaia BH1 and BH2 (El-Badry et al., 2023a, b). For example, Tanikawa et al. (2024a) predict that the formation efficiency per unit mass of dynamically formed BH + star binaries increases by less than a factor of $\sim 5$ between $\rm[M/H]=0$ and $\rm[M/H]=-2$.
@@ acknowledgments
I thank Jim Fuller, Re’em Sari, Jared Goldberg, Thomas Tauris, Hans-Walter Rix, Vedant Chandra, and Rene Andrae for conversations that were influential to this work, and Jakub Klencki for making his MESA inlists publicly available. This research was supported by NSF grant AST-2307232.

### v2

@@ 4.1 Wide BH companions are overrepresented at low metallicity
Figure 5: Metallicity distribution of red giants within 1 kpc of the Sun, calculated from the Gaia XP metallicity catalog of Andrae et al. (2023, black) and APOGEE DR17 (Abdurro’uf et al., 2022, red). The metallicity of Gaia BH3 is marked with a shaded region: $\rm[M/H]=-1.82$ according to Andrae et al. (2023), and $\rm[M/H]=-2.2$ according to Gaia Collaboration et al. (2024). The fraction of giants with metallicities as low as Gaia BH3 is only $\sim 0.0001$.
@@
Both the XP and APOGEE samples suggest that only 1 in $\sim 10^{4}$ giants within 1 kpc have $\rm[M/H]<-2.0$. This very small fraction strongly implies that wide BH companions are overrepresented at low metallicity. If they were not, we could expect to find $\sim 10^{4}$ BH companions – just to red giants – within 1 kpc. The true number of BH + giant binaries within 1 kpc is unknown, but 0 have been discovered so far: the nearest such system is Gaia BH2 at $d=1.16$ kpc (El-Badry et al., 2023b). It seems quite unlikely that $10^{4}$ remain to be discovered (although published Gaia data are not yet sensitive to orbital periods as long as that of Gaia BH3), and so a more plausible explanation is that wide BH companions are more common at low metallicity. Early results from searches for neutron star companions in the Gaia data similarly imply that they are overrepresented at low metallicity (El-Badry et al., 2024a). More work is needed to understand why.
@@ 4.2 More low-metallicity BHs are expected in DR4
There is also reason to believe that more BHs orbited by metal-poor stars exist in the solar neighborhood. Such stars are uniformly $\gtrsim 12$ Gyr old. This means that giants sample a narrow mass range of roughly $(0.78-0.80)\,M_{\odot}$. For a Kroupa IMF, there are $\sim$165 stars with $M=(0.1-0.78)\,M_{\odot}$ for every star with $M=(0.78-0.80)\,M_{\odot}$, suggesting that even within 1 kpc, other massive BHs are likely orbited by lower-mass metal-poor stars that are still on the main sequence. The mass distribution of companions to BHs of course may be different from the IMF, but it seems unlikely that $0.8\,M_{\odot}$ companions would be significantly more common than companions with mass of $0.7\,M_{\odot}$ or $0.6\,M_{\odot}$. At the distance and extinction of Gaia BH3, such companions would have $G<14$ for $M>0.75\,M_{\odot}$, or $G<16$ for $M>0.61\,M_{\odot}$. These magnitudes are too faint for Gaia to have measured multi-epoch RVs from RVS spectra, but bright enough that the astrometric orbit would be resolved with $\rm SNR\gtrsim 100$. With careful processing of the astrometric data, a source like Gaia BH3 could likely be detected even at $G=18-19$ – where individual astrometric measurements have uncertainties of 1-3 mas in the along-scan direction, a factor of 10-30 smaller than Gaia BH3’s photocenter ellipse (Lindegren et al., 2018, 2021; Holl et al., 2023) – although spurious solutions will be more abundant at low SNR. Gaia BH3-like systems can thus quite plausibly be detected with significantly lower-mass secondaries, and to larger distances, than Gaia BH3 itself.
@@ 5.1 Formation from an isolated binary
I have shown that wide BH + stellar binaries with orbits similar to Gaia BH3 can in principle be produced through isolated binary evolution, without the BH progenitor ever interacting with its companion. However, the orbit of Gaia BH3 in particular appears unlikely to have formed from this channel if we accept that the binary was born in the progenitor of the ED-2 stream and remained bound to it when the BH formed: matching the system’s observed period and eccentricity while accommodating the BH progenitor as a red supergiant requires a natal kick of velocity $\gtrsim 20\,{\rm km\,s^{-1}}$, and such a kick would have ejected the binary from a low-mass cluster. Given the old age of the systems and the short lifetime of the BH progenitor, it seems unlikely that the binary was ejected from the stream with a significant kick and still remains closely associated with the stream in phase space
The main uncertainty in these calculations is in the maximum radii predicted by the evolutionary models. For example, Figure 1 shows that rapidly-rotating BH progenitors could have undergone CHE and never overflowed their Roche lobes in any plausible pre-SN orbit of Gaia BH3. The radii of red supergiants are also sensitive to a variety of uncertain inputs, such as the convective mixing length and the treatment of overshooting and semiconvection (Chun et al., 2018; Schootemeijer et al., 2019; Goldberg & Bildsten, 2020; Klencki et al., 2021; Romagnolo et al., 2023). Indeed, Iorio et al. (2024) perform similar calculations to those presented here but employ evolutionary tracks from the MIST project (Choi et al., 2016), which predict maximum radii of only $\sim 200\,R_{\odot}$. Based on these models, they find that Gaia BH3’s properties are consistent with formation from an isolated binary. The modeling choices adopted here produce red supergiant radii in reasonably good agreement with observations of massive stars in the LMC and SMC (e.g. Schootemeijer et al., 2019), which are the lowest-metallicity galaxies currently observable with a large population of evolved massive stars. The maximum radii they predict at $\rm[M/H]=-2.2$ are not significantly different at fixed mass from those predicated at $\rm[M/H]=-0.7$, the metallicity of the SMC. No one has ever observed a massive star with a metallicity as low as Gaia BH3, so the evolutionary models are untested in the lowest-metallicity regime.
@@ 5.2 Dynamical formation
With formation from a primordial binary unlikely, the alternative is dynamical assembly in the cluster that formed ED-2. Cluster formation models have also previously been found promising for forming the other two BH binaries discovered so far from Gaia data (Rastello et al., 2023; Di Carlo et al., 2023; Tanikawa et al., 2024b), and for BH binaries found in globular clusters with spectroscopic surveys, some of which have similar parameters to the Gaia BHs (Kremer et al., 2018; Gieles et al., 2021).
Recent modeling by Marín Pina et al. (2024) predicts a population of binaries with properties similar to Gaia BH3 that escape from Monte Carlo cluster models. They predict a separation distribution for dynamically-formed binaries that peaks at close separations but is relatively flat in log period over $P_{\rm orb}=10^{0-4}$ d. All the clusters in the library analyzed by Marín Pina et al. (2024) have total masses significantly larger than Balbinot et al. (2024) estimate for the progenitor of the ED-2 stream, so it would be useful to model the formation of Gaia BH3-like binaries in lower-mass clusters. The most extensive effort in this direction thus far was undertaken by Tanikawa et al. (2024b, a), who also predict a period distribution that is relatively flat in log space for clusters with $M_{\rm cluster}=10^{3}\,M_{\odot}$. A full exploration of the low-mass cluster regime is still warranted.
Marín Pina et al. (2024) predict that the formation efficiency of wide BH binaries per unit mass increases as $\sim M_{\rm cluster}^{-1.2}$, such that low-mass clusters will contribute the large majority of BH+star binaries and the predicted number of detectable binaries depends sensitively on the behavior of the adopted cluster mass function at the low-mass end. Integrating down to $M_{\rm cluster}=10^{2}\,M_{\odot}$ (which may be optimistic, since the lowest-mass clusters may not form any BHs or be dense enough for dynamical interactions to be important), they predict $\lesssim 1$ low-metallicity BH binary within 1 kpc. The empirical rate considerations in Section 4.2 suggests that the true number of such binaries may be larger than this, but the Poisson uncertainty associated with $N=1$ detection still allows for the possibility that Gaia BH3 was a lucky discovery.
5.3 The possibility of pollution
It may be possible to discriminate between dynamical and isolated binary evolution channels for future BH discoveries on the basis of the abundances of the companion stars. Gaia Collaboration et al. (2024) and Balbinot et al. (2024) show that Gaia BH3’s companion does not show any significant abundance anomalies, and they suggest that this favors a dynamical formation channel. In support of this possibility, some of the neutron star binary candidates identified by El-Badry et al. (2024a) do show abundance anomalies, though their origin is still uncertain.
For Gaia BH3, I find that pollution from SN ejecta would likely not be detectable, even if the system had formed from isolated binary evolution. For a plausible separation of 10 au and companion star radius of $1\,R_{\odot}$ at the time of the SN, the companion would have subtended a fraction $f\sim 5\times 10^{-8}$ of the sky as viewed from the BH progenitor. Assuming $20\,M_{\odot}$ of isotropic ejecta, only $\sim 10^{-6}\,M_{\odot}$ of ejecta – likely consisting mostly of oxygen (Thielemann et al., 1996) – would have been deposited on the star. Since the star is now a giant, its convective zone extends deep into the interior: from a MESA model, I find that the outer $\sim 0.35\,M_{\odot}$ is likely convective. If $\sim 7\times 10^{-7}\,M_{\odot}$ of oxygen had been accreted, the ejecta would have been mixed with $\sim 3\times 10^{-5}\,M_{\odot}$ of oxygen already in the envelope, changing the star’s surface abundances by at most a few percent, which is unlikely to be detectable. Chances of detecting pollution would be much higher if the companion star were still on the main sequence, where the outer convective zone is predicted to contain $<10^{-3}\,M_{\odot}$.
5.4 Future prospects
@@
Gaia BH3 is just at the separation beyond which BH+star binaries can avoid a common envelope. Irrespective of whether Gaia BH3 in particular formed from an isolated binary, the models presented here imply that isolated binary evolution produces binaries with orbits similar to Gaia BH3, and future Gaia data releases are likely to find these systems. This makes the system different from Gaia BH1 and BH2, which most studies have found difficult to produce through isolated binary evolution under standard assumptions (El-Badry et al. 2023a; Rastello et al. 2023; Tanikawa et al. 2024b, though see Kotko et al. 2024 for a different view).
@@
Constraints on the separation distribution of BH binaries from DR4 will provide further insights into their dominant formation channel. If these systems form mainly by isolated binary evolution, models predict the BH + low mass star period distribution to rise steeply at $P_{\rm orb}\gtrsim 10$ yr (e.g. Breivik et al., 2017; Chawla et al., 2022), the period beyond which a significant fraction of binaries will have avoided a common envelope inspiral. For dynamical formation channels, the period distribution will depend on the mass of the clusters in which most systems are formed, but there is little reason to expect a steep increase at $P_{\rm orb}\gtrsim 10$ yr, and models published so far predict a period distribution that is flat or declines at longer periods (Tanikawa et al., 2024b; Marín Pina et al., 2024). Fortunately, Gaia will obtain an 11-year observational baseline (Gaia Collaboration et al., 2016), enabling robust orbital constraints at periods up to $\sim 20$ years, and perhaps even significantly longer (e.g. Andrews et al., 2023).
@@
The metallicity distribution of BH binaries may also help constrain how they formed. The vast majority of clusters that have contributed stars to the solar neighborhood had metallicities near solar. If dynamical interactions form most wide BH binaries, most systems should likely have metallicities near solar, similar to Gaia BH1 and BH2 (El-Badry et al., 2023a, b). For example, Tanikawa et al. (2024a) predict that the formation efficiency per unit mass of dynamically formed BH + star binaries increases by less than a factor of $\sim 5$ between $\rm[M/H]=0$ and $\rm[M/H]=-2$. Curiously, the empirical data so far suggests a significant increase in the prevalence of both BH and neutron star companions at low metallicity (El-Badry et al., 2024b, a).
@@ acknowledgments
I thank the anonymous referee for comments that significantly improved the paper. I am also grateful to Jim Fuller, Re’em Sari, Jared Goldberg, Thomas Tauris, Hans-Walter Rix, Vedant Chandra, and Rene Andrae for conversations that were influential to this work, and Jakub Klencki for making his MESA inlists publicly available. This research was supported by NSF grant AST-2307232.
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.